When a voltage of 100 volts at 50 Hz is applied to a series connection of a resistor A

and an inductor A, the current taken is 8 A and the power is 120 watts. When applied

to another series connection of a resistor B and an inductor B, the current is 10 A and

the power is 500 watts. What current and power will be taken when 100 V is applied

to the two coils in series?


1
Expert's answer
2021-03-01T04:57:51-0500

"V=100V\newline f=50Hz"


CaseA:

"Ia=8A\newline Pa=120W"

"Ra=Pa/Ia^2=120/8^2=1.875\Omega"

"Za=V/Ia=100/8=12.5\Omega"

"Xa=\sqrt{Za^2-Ra^2} =\sqrt{12.5^2-1.875^2} =12.36\Omega"

"\implies La\omega=12.36, La=12.36/2\pi50=39.34mH"

CaseB:

"Ib=10A\newline Pb=500W"

"Rb=Pb/Ib^2=500/10^2=5\Omega"

"Zb=V/Ib=100/10=10\Omega"

"Xb=\sqrt{Zb^2-Rb^2} =\sqrt{10^2-5^2} =8.66\Omega"

"\implies Lb\omega=8.66, Lb=8.66/2\pi50=27.56mH"


When La and Lb are connected in series current "Iab=V\/(Xa+Xb) =100\/(12.36+8.66) =4.76A"

Power= 0 , as both are inductors there is no power loss.